OpenStudy (anonymous):

Solve. x^2 – 10x = –13 5 ± 2 sqrt 3 5 ± sqrt 38 5 ± 2 sqrt 2 5 ± 3 sqrt 2

OpenStudy (anonymous):

first move everything to one side x^2 - 10x + 13 use the quadratic formula. (-b +/- sqrt(b^2 + 2ab))/2a

OpenStudy (anonymous):

oops i mean - 4ac for the quadratic. you'll get (-10 +/- sqrt(100-52))/2 so -5 +/- sqrt (12) someone correct me if i'm wrong please.

OpenStudy (anonymous):

\[\frac{10\pm\sqrt{10^2-4(1)(13)}}{2(1)}\] \[\frac{10\pm\sqrt{100-52}}{2}\] \[\frac{10\pm\sqrt{48}}{2}\] \[\frac{10\pm\sqrt{3(16)}}{2}\] \[\frac{10\pm4\sqrt{3}}{2}=5\pm2\sqrt{3}\]

OpenStudy (anonymous):

The quadratic setup was right. You had until you started reducing.

OpenStudy (anonymous):

you have the answer but i would like to point out that if you see the middle term is even and you 'complete the square' you can avoid a lot of reducing

OpenStudy (anonymous):

\[x^2-10x-13\] \[(x-5)^2=-13+25=12\] \[x-5=\pm2\sqrt{3}\] etc

OpenStudy (anonymous):

formula forces a denominator on you (2a) which is unnecessary if the coefficient of the "middle term" is even (assuming it is monic)